OpenStudy (countrygirl1431):

Make the following conversion. 34.9 cL = _____ hL 0.349 0.0349 0.00349 349,000

OpenStudy (countrygirl1431):

@LynFran

OpenStudy (lynfran):

1 cL=1* 0.0001 hL so 34.9 cl= 34.9* 0.0001hL

OpenStudy (countrygirl1431):

0.00349?

OpenStudy (lynfran):

correct

OpenStudy (countrygirl1431):

thanks

OpenStudy (lynfran):

welcome